This Quiz: 8pts Two standardized tests, test A and test B, use very dfferent scaes Assume byN 1500, 125) and scores on teset 8 can be modeied by N(17, 2) if an applicent to a univernsity he has taken sest B and scored 24, compare these studenis scores using z-vaiues Which one has a higher relative score? Explain The z-value ofthe test A sore isZA.□ Round to two decimal places as needed) The 2-value of the test B score is 7 Round to two decimai places as needed.) O A. The test B score is the higher relative score because it is farther above the mean in standand O B. The test A score is the higher relative score because it is higher than the score on test O C. The test A score is the higher relative score because it is tarther above the mean in standand deviation units than the test 8score OD, The test B soore is the higherrelative score because it higher tan the score on test A deviation units than the test A scoe
Answer #1

z value for test A score = (1900 - 1500)/125 = 3.2

z value for test B score = (24 - 17)/2 = 3.5

Option A is correct.